Graph each number on the complex plane and find its absolute value..
8. 3 - 4i
9. 8 + 10i
10. 9 + 3i

Answer by stanbon(75887)

Graph each number on the complex plane and find its absolute value..
8. 3 - 4i
Plot the point (3,-4); abs. value = sqrt(3^2 + 4^2) = 5
-----------------------
9. 8 + 10i
Plot the point (8,10); abs. value = sqrt(8^2+10^2) = 12.806...
-----------------------
10. 9 + 3i
Plot the point (9,3); abs. value = sqrt(9^2+3^2) = 9.4868...
====================
Cheers,
Stan H.